OpenStudy (xapproachesinfinity):

suppose f is function with property \[\large |f(x)|\leq x^2\] for any x in R a) show that f(0)=0 b) show that f'(0)=0

OpenStudy (xapproachesinfinity):

so far no idea hehe :)

OpenStudy (kainui):

Here's my guess, but I'm not entirely sure it's allowed since I think the square has to be under the square root, not inside out like this: \[\Large \sqrt{f(x)}^2 \le x^2 \\ \Large \pm \sqrt{f(x)}\le x\]

jimthompson5910 (jim_thompson5910):

|f(x)| <= x^2 turns into -x^2 <= f(x) <= x^2 then you plug in x = 0 -x^2 <= f(x) <= x^2 -0^2 <= f(0) <= 0^2 0 <= f(0) <= 0 so f(0) = 0 I'm not sure how to do part b

OpenStudy (xapproachesinfinity):

sounds good to me @Kainui

OpenStudy (freckles):

differentiate !

OpenStudy (freckles):

your inequality

OpenStudy (freckles):

-2x<f'(x)<2x

OpenStudy (xapproachesinfinity):

damn @jim_thompson5910 was thinking to do that thought it wouldn't work awesome!!

OpenStudy (xapproachesinfinity):

is that allowable? @freckles

jimthompson5910 (jim_thompson5910):

freckles, you forgot about the absolute value

OpenStudy (freckles):

well that inequality is true for x>0 and then 2x<f'(x)<-2x for x<0

jimthompson5910 (jim_thompson5910):

if y = |x|, then dy/dx = |x|/x or dy/dx = x/|x|

OpenStudy (freckles):

\[-x^2 \le f(x) \le x^2 \] so we can't just differentiate everything here?

OpenStudy (xapproachesinfinity):

hmm doing cases yes that good! both way we get f'(0)=0

jimthompson5910 (jim_thompson5910):

oh true, just break it up and then differentiate, nvm then

OpenStudy (xapproachesinfinity):

yeah makes sense to me, awesome guys ^^